OpenStudy (pumpkin1102):

A student is trying to solve the system of two equations given below: Equation P: x + y = 8 Equation Q: 2x + 5y = 24 Which of the following steps can be used to eliminate the x term? 2(x + y = 8) −2(x + y = 8) −1(2x + 5y = 24) −2(2x + 5y = 24)

OpenStudy (pumpkin1102):

it's not A

Mehek (mehek14):

If it's not A, there's only 1 other option that can work

OpenStudy (anonymous):

the answer is B....the others are eliminated by substitution.

OpenStudy (pumpkin1102):

why

Mehek (mehek14):

the other equation doesn't need to be multiplied by anything you need to multiply the first one to eliminate x

OpenStudy (pumpkin1102):

?

OpenStudy (anonymous):

The others will cancel out both x and Y. (The correct way is substituting the value of x from the first eqn into the second. and thus eliminating x.. So..x+y=8.....x=8-y put this into second equation 2(8-y)+5y=24

OpenStudy (pumpkin1102):

i think that just confused me more gosh who invented school

OpenStudy (phi):

Equation P: x + y = 8 Equation Q: 2x + 5y = 24 if you want the x's to "go away" then you want them to add up to 0 right now you have 1 x and another 2x's and they add up to 3 x Not what you want. But look at the 2x in equation Q

OpenStudy (phi):

what do you want to do to the 2x ? you want to take away 2x how do we takeaway 2x ? we subtract: 2x - 2x or add -2x 2x+ -2x

OpenStudy (pumpkin1102):

so you want to get rid of 2x by -2

OpenStudy (phi):

yes, the idea is to multiply by -2 which equation? the equation that has only x, because -2 * x gives -2x and that will add to the 2x in the other equation and give zero.

OpenStudy (pumpkin1102):

oh I think I actually get it

OpenStudy (anonymous):

It wants us to eliminate x, and not make the equation =0 okay...second answer becomes= -2x-2y=-16 and when adding this into the second equation...x is cancelled out.and 3y=8 y=8/3

OpenStudy (pumpkin1102):

I would medal someone but I don't know who to give it to

OpenStudy (pumpkin1102):

help me decide who to medal I don't want to leave anyone out

OpenStudy (pumpkin1102):

there @magicalsd

OpenStudy (anonymous):

give it to the other person...he/she might need it more...i am used to working hard

OpenStudy (pumpkin1102):

are u sure?

OpenStudy (anonymous):

can u give to both of us?

OpenStudy (pumpkin1102):

no the system won't let me just keep it he/she left anyways

OpenStudy (anonymous):

thanks a lot and good luck to you.Hope you succeed.